OpenStudy (anonymous):

Factor the following polynomial: x2 − 8x + 12 (x − 4)(x − 3) (x − 2)(x − 6) (x − 2)(x + 6) (x + 4)(x − 2)

OpenStudy (anonymous):

x^2 − 8x + 12 = x^2 -2x -6x + 12 = x(x-2) -6(x-2) = (x-6)(x-2)